OpenStudy (anonymous):

A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 800 centimeters tall. I got Un=800+.05(n-1) is this correct?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

No sorry, it is 0.5, not 0.05.

OpenStudy (anonymous):

great thanks!!!

OpenStudy (turingtest):

you wrote a recursive function, it looks like, much fancier than what they seem to ask for is that supposed to be \[u_n=800+0.5u_{n-1}\]or\[u_n=800+0.5u(n-1)\]?

OpenStudy (anonymous):

2nd one

OpenStudy (turingtest):

then it's not quite right, because when the tree arrives we would call that week 0, and if you plug n=0 you get less than 800 cm

OpenStudy (turingtest):

actually the problem is not well-worded about that, but the safer way to go is to start counting from wek zero

OpenStudy (turingtest):

week*

OpenStudy (anonymous):

I think it is just an arithmetic sequence question where it wants you to see that 800 cm is the first term (a) and the common difference is 0.5 cm (d).

OpenStudy (turingtest):

the wording is not the best, but it does say "function" not "construct an arithmetic sequence" and this would not be an arithmetic seuence if you don't have n in the subscript

OpenStudy (turingtest):

function of the height in terms of weeks after arrival is height(week)=initial height + (growth/week)X(weeks after arrival)

OpenStudy (turingtest):

this seems to me the most straightforward way of reading the problem, but again, it should specify "On week zero, the height is..." to clarify

OpenStudy (anonymous):

I know, but I've been answering a few questions from the same person and they were all about finding the difference and first term of an arithmetic sequence so that's what I assumed this question was about again.

OpenStudy (turingtest):

fair enough, I'll trust your experience with this particular series of questions over the ambiguity of the wording